FIND THE VALUE OF x. PLEASE GIVE ME DETAILED SOLUTION WITH EACH AND EVERY STEP .

FIND  THE  VALUE  OF  x.
PLEASE  GIVE  ME  DETAILED  SOLUTION  WITH  EACH  AND  EVERY  STEP .

Grade:10

1 Answers

Arun
5 years ago
1/(2a+b+2x) - 1/(2a) = 1/(2x) + 1/(b)
 
[2a - (2a+b+2x)] / [2a(2a+b+2x)] = (b+2x) / [2bx]
 
-(2x+b) / [2a(2a+b+2x)] = (b+2x)/ (2bx)
 
-1 / [a(2a+b+2x)] = 1/(bx)
 
-bx = 2a2 + ab + 2ax
 
So, x = [a(2a+b)] / [-(b + 2a)]
 
 So, x = -a & -b/2